Selected current year-end financial statements of Cabot Corporation follow. (All sales were on credit; selected balance sheetReq 10 Req 1 and 2 Req 5 Req 7 Req 9 Req 3 Req 4 Req 6 Req 8 Req 11 Compute the current ratio and acid-test ratio. (1) CurrenReq 5 Req 6 Req 1 and 2 Req 4 Req 3 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the days sales uncollected. Days Sales UncollecReq 9 Req 1 and 2 Req 4 Req 6 Req 7 Req 5 Req 10 Req 11 Req 3 Req 8 Compute the inventory turnover. Inventory Turnover (4) InReq 1 and 2 Req 5 Req 3 Req 6 Req 7 Req 9 Req 11 Req 4 Req 8 Req 10 Compute the days sales in inventory. Days Sales in InveReq 6 Req 1 and 2 Req 11 Req 5 Req 7 Reg 3 Req 4 Req 8 Req 9 Req 10 Compute the debt-to-equity ratio. Debt-to-Equity Ratio (6Reg 5 Req 1 and 2 Req 11 Req 3 Req 4 Req 6 Req 7 Req 8 Req 9 Req 10 Compute the times interest earned. (7) Times Interest EarReq 1 and 2 Req 3 Req 9 Req 4 Req 8 Req 5 Req 6 Req 7 Req 10 Req 11 Compute the profit margin ratio. Profit Margin Ratio (8)Req 1 and 2 Req 6 Req 8 Req 3 Req 4 Req 9 Req 10 Req 5 Req 7 Req 11 Compute the total asset turnover. Total Asset Turnover (9Req 9 Req 6 Req 8 Req 1 and 2 Req 10 Req 11 Req 3 Req 4 Req 5 Req 7 Compute the return on total assets. Return on Total AssetReq 6 Req 1 and 2 Req 7 Reg 3 Req 4 Req 5 Req 8 Req 9 Req 10 Req 11 Compute the return on common stockholders equity. Return
Answer #1

First 5 parts are being answered here.

1. Current ratio = Current assets / Current liabilities

Current assets = Cash + short term investments + accounts receivable + merchandise inventory + Prepaid expenses

Current assets = $10000 + $9400 + $30400 + $38150 + $2450 = $90400

Current liabilities = Accounts payable + Accrued wages payable + Income taxes payable

Current liabilities = $15500 + $3800 + $3200 = $22500

Now, putting these values in the current ratio formula, we get,

Current ratio = $90400 / $22500 = 4.02

2. Acid test ratio = Current assets – Merchandise inventory – Prepaid expenses / Current liabilities

Current assets = $90400, Current liabilities = $22500, Inventories = $38150, Prepaid expenses = $2450

Acid test ratio = ($90400 - $38150 - $2450) / $22500

Acid test ratio = $49800 / $22500 = 2.21

3. For day's sales uncollected, we need to calculate accounts receivable turnover ratio first.

Accounts receivable turnover ratio = Credit sales / Accounts receivables

Credit sales = $454600, Accounts receivable = $30400

Putting these values in the accounts receivable turnover ratio formula, we get,

Accounts receivables turnover ratio = $454600 / $30400 = 14.9539473684 times

Now, we will calculate days's sales uncollected by the following formula,

Days' sales uncollected = 365 / Accounts receivable turnover ratio

Accounts receivable turnover ratio = 14.9539473684 times ( as computed above)

Putting these values in the Days' sales in inventory ratio formula, we get,

Days' sales uncollected  = 365 / 14.9539473684 = 24.41 days

4. Inventory turnover ratio = Cost of the goods sold / Average inventory

where, Average inventory = Beginning inventory + Ending inventory / 2

For current year:

Ending inventory = $38150, Beginning inventory = Ending inventory of prior year = $48900

Average inventory = ($38150 + $48900) / 2 = $87050 / 2 = $43525

Cost of the goods sold = $297650

Putting these values in the inventory turnover ratio formula, we get,

Inventory turnover ratio = $297650 / $43525 = 6.84 times

(5) Days' sales in inventory = 365 / Inventory turnover ratio

Inventory turnover ratio = 6.84 times ( as computed above)

Putting these values in the Days' sales in inventory ratio formula, we get,

Days' sales in inventory ratio = 365 / 6.84 = 53.36 days